### v1

# Training-Free ANN-to-SNN Conversion for High-Performance Spiking Transformer

Abstract
Leveraging the event-driven paradigm, Spiking Neural Networks (SNNs) offer a promising approach for constructing energy-efficient Transformer architectures. Compared to directly trained Spiking Transformers, ANN-to-SNN conversion methods bypass the high training costs. However, existing methods still suffer from notable limitations, failing to effectively handle nonlinear operations in Transformer architectures and requiring additional fine-tuning processes for pre-trained ANNs. To address these issues, we propose a high-performance and training-free ANN-to-SNN conversion framework tailored for Transformer architectures. Specifically, we introduce a Multi-basis Exponential Decay (MBE) neuron, which employs an exponential decay strategy and multi-basis encoding method to efficiently approximate various nonlinear operations. It removes the requirement for weight modifications in pre-trained ANNs. Extensive experiments across diverse tasks (CV, NLU, NLG) and mainstream Transformer architectures (ViT, RoBERTa, GPT-2) demonstrate that our method achieves near-lossless conversion accuracy with significantly lower latency. This provides a promising pathway for the efficient and scalable deployment of Spiking Transformers in real-world applications.
Index Terms:
Spiking Neural Networks, ANN-to-SNN Conversion, Spiking Transformer, Few Spikes Neuron.
I Introduction
Spiking Neural Networks (SNNs) have garnered widespread attention due to their sparse spike-driven computing paradigm [1, 2]. Unlike traditional Artificial Neural Networks (ANNs) transmit information via continuous-valued activations, SNNs employ sparse binary spikes as information carriers, thereby offering superior energy efficiency for resource-limited devices [3, 4]. Recently, the SNNs community has incorporated high-performance Transformer [5] architectures into SNNs to leverage both the powerful representation capabilities of Transformer and the inherent energy efficiency of the SNNs [6, 7, 8]. These methods provide significant performance improvements and broaden the application scenarios of SNNs.
To obtain Transformer-based SNNs, two approaches exist: direct training (DT) [9, 10] and ANN-to-SNN (A2S) conversion [11, 12]. DT employs surrogate gradients [13] and backpropagation through time (BPTT), but exhibits inaccurate gradient approximation and $\mathcal{T}\times$ training overhead. In contrast, A2S converts pre-trained ANNs to SNNs, maintaining inference efficiency while avoiding training overhead. A2S methods [14, 15] achieve lossless conversion in convolutional structures by establishing an equivalence between RELU activations and firing rates [16, 17]. However, Transformer conversion remains challenging due to complex nonlinear operations.
Nonlinear operations in Transformer architectures can be divided into two categories: single-variable operations (e.g., GELU, Tanh) and multi-variable operations (e.g., variable-variable floating-point (FP) multiplications, LayerNorm). These operations lead to notable performance degradation [17]. SpikeZIP-TF [18] replaces single-variable functions and applies fine-tuning to adjust pre-trained ANN weights, though at the cost of additional training overhead. STA [19] approximates all nonlinear functions using group operators, but this approach incurs increased inference latency. Therefore, it is necessary to design a high-performance conversion that balances training cost and inference efficiency.
Inspired by the temporal coding mechanisms observed in biological neurons, Few-spikes (FS) neurons [20] are proposed. Unlike rate-coded approaches, it leverages spike timing and patterns to represent neuronal activation states. This characteristic enables near-lossless fitting for single-variable operations. However, vanilla FS neurons suffer from significant performance degradation in large-scale deep networks. Moreover, their design based on single-input fitting limits their ability to handle the multi-variable operations in Transformer architectures.
In this paper, we propose a training-free A2S conversion framework tailored for Transformer architectures, enabling near-lossless conversion without modifying the weights of pre-trained ANNs. We begin by theoretically and empirically analyzing the key challenges associated with FS neuron-based conversion, particularly in approximating nonlinear operations. To overcome these challenges, we introduce a novel Multi-Basis Exponential Decay (MBE) neuron, which effectively approximates various nonlinear operations. Built upon MBE neurons, our method achieves both low inference latency and near-lossless Transformer conversion. The main contributions of this paper are as follows:
  1. 1.
    We systematically analyze the incompatibility between FS neurons and Transformer architectures, focusing on excessive dependence on initialization (EDI) and global suboptimality (GSO) problems, which hinder convergence and degrade conversion performance.
  2. 2.
    We introduce a MBE neuron with exponential decay strategy and multi-basis encoding method. The decay strategy enables multi-resolution representations, while multi-basis encoding enhances the near-lossless approximation of diverse nonlinear operations.
  3. 3.
    We propose an A2S framework based on MBE neurons that efficiently approximates various nonlinear operations in Transformer architectures, including variable-variable FP multiplications, GELU, Softmax, and LayerNorm. It achieves near-lossless conversion with fast inference and require no training on source ANNs.
  4. 4.
    Extensive experiments on diverse tasks (CV, NLU, NLG) and architectures (ViT, RoBERTa, GPT-2) demonstrate that our method achieves near-lossless conversion with reduced latency, yielding competitive results among existing Transformer-based A2S methods.
II Related work
Existing conversion paradigms can be categorized into two types based on whether extra training on the source ANNs is required: training-dependent and training-free. Training-dependent conversion typically requires additional ANNs training based on theoretical ANN-SNN equivalence. Clip-Floor-Shift activations [16] replace ReLU to better match spiking neuron behavior. A two-phase training [21] first optimizes ANN weights, then adjusts neuron thresholds, while activation range constraints and membrane potential initialization [22] help further reduce conversion error. [23] introduce a two-stage approach to handle quantization, pruning, and residual errors. [24] propose a unified framework treating spike-rate mapping as a differentiable problem. For Transformer architectures, a QANN is integrated with spatiotemporal properties of spiking neurons for lossless conversion [18]. While effective, these methods introduce computational overhead through additional source ANNs training.
Training-free conversion directly transforms pre-trained ANNs via structure and parameter reuse without fine-tuning. Early works align SNN thresholds with ANN activations via threshold balancing [11, 12], but longer timesteps are required. Signed spiking neurons [25] support dynamic vision data. In [19], training-free Transformer conversion is achieved using universal group operators and spatial rectification self-attention. SpikedAttention [26] enables spike-driven Transformer A2S via trace-based matrix multiplication and winner-take-all spike shifting, yet retains non-spiking LayerNorm. ECMT [27] reduces latency in Transformer SNNs, yet still relies on floating-point operations. While retraining is avoided, these methods often require long timesteps or retain non-spiking components, limiting SNNs’ energy efficiency.
III Preliminary & Problem Analysis
III-A Preliminary
Inspired by authentic electrophysiological characteristics of human brain neurons, FS neuron is proposed to address the inherent trade-off between spike count and accuracy in A2S [20]. By introducing learnable parameters including neuron threshold, membrane potential reset value, and spike intensity, FS neuron approximates ANN activation functions with very few spikes. Mathematically, the membrane potential of FS neuron can be computed as:
$u[t+1]=u[t]-r[t]\cdot s[t],$ (1)
where $t\in[0,T-1]$ denotes timesteps, $u[t]$ represents the membrane potential, $s[t]\in\{0,1\}$ is the binary spike, and $r[t]$ is the learned reset value. The membrane potential has no decay, and its initial value is typically set to the input from the previous layer, i.e., $u[0]=x$. When the membrane potential exceeds a learned threshold, a spike is generated, which is described as follows:
$s[t]=\mathcal{H}\left(\left(x-\sum\nolimits_{t^{{}^{\prime}}=0}^{t-1}r[t^{{}^{\prime}}]\cdot s[t^{{}^{\prime}}]\right)-V_{th}[t]\right),$ (2)
where $\mathcal{H}(\cdot)$ is the Heaviside step function and $V_{th}[t]$ is the learned firing threshold. After each spike emission, the reset mechanism is invoked to update the membrane potential, as described in Eq.(1). Noteworthy, the spike $s[t]$ emitted by FS neuron at time $t$ is amplified by learned spike intensity $d[t]$. After $T$ timesteps, the neuron integrates the weighted spike train and forwards it to the next layer:
$\hat{f}(x)=\sum\nolimits_{t=0}^{T-1}d[t]s[t]\approx f{(x)},$ (3)
where $\hat{f}(x)$ is the approximation value produced by FS neuron and $f(x)$ is the output of activation function in ANNs.
III-B Problem Analysis
We conduct a systematic analysis of FS neurons to understand their limitations in Transformer-based A2S tasks.
Figure 1: Loss landscape of FS neurons in approximating the ReLU activation function. The 3D landscape of $\mathcal{L}_{Binary}$ and $\mathcal{L}_{Random}$ from two different initialization.
Approximation Error Analysis To qualitatively analyze how the FS neuron design impacts performance, following [19], we consider errors from three sources: insufficient sampling, limited parameterization, and spiking quantization. This yields the following error bound:
Theorem 1 (FS Neuron Error Bound).
Let $f:[a,b]\rightarrow\mathbb{R}$ be a target activation function, and let $\hat{f}_{T}^{(M)}(x)$ denote the output of an FS neuron with $T$ timesteps trained on $M$ samples. Then the total approximation error satisfies:
$\varepsilon\leq\mathcal{O}\Biggl{(}\underbrace{\sqrt{\frac{T\log T\log M}{M}}}_{\text{Empirical Gap}}+\underbrace{\frac{\mathcal{L}_{f}|y|_{\max}}{T}}_{\text{Parametric Gap}}+\!\underbrace{\frac{\|d\|_{1}}{T}}_{\text{Quantization Gap}}\Biggr{)},$
where $\varepsilon=\mathbb{E}[|f(x)-\hat{f}_{T}^{(M)}(x)|]$, $\mathcal{L}_{f}$ is the Lipschitz constant of $f$, $|y|_{\max}$ the maximum value of $f(x)$, and $\|d\|_{1}$ is the $L_{1}$-norm of spike intensities. Proof in Appendix A.
III-B1 Excessive Dependence on Initialization
From Theorem 1, the quantization gap $\frac{\|d\|_{1}}{T}$ depends on spike intensities $d[t]$, which are determined through optimization from initialization. Poor initialization may produce suboptimal $d[t]$ that increase $\|d\|_{1}$, thereby enlarging the approximation error. We conduct experiments to validate this: random initialization yields $\text{MSE}_{\text{random}}=1.5\times 10^{-3}$, while binary initialization ($V_{\mathrm{th}}[t]=r[t]=d[t]=2^{T-t}$) achieves $\text{MSE}_{\text{binary}}=9.4\times 10^{-5}$, representing a 93.29% performance degradation under random initialization. As shown in Fig. 1, binary initialization results in a smoother and more stable optimization landscape, while random initialization leads to multiple local minima. This further corroborates the sensitivity of FS neurons to initialization quality.
Figure 2: Critical interval approximation failure of nonlinear operations in Transformer using FS neurons.
III-B2 Global Suboptimality
GSO stems from the parametric gap term $\frac{\mathcal{L}_{f}|y|_{\max}}{T}$ in Theorem 1, revealing FS neurons’ limitations when approximating functions with non-uniform complexity. Transformer nonlinearity inputs concentrate near zero (Statistical analysis in Appendix C), where activation functions such as GELU and SiLU have high curvature and large local Lipschitz constants $\mathcal{L}_{f}^{\text{local}}\gg\mathcal{L}_{f}$. This creates an amplified local gap $\varepsilon_{\text{param}}^{\text{local}}=\mathcal{O}(\mathcal{L}_{f}^{\text{local}}|y|_{\max}^{local}/T)$, which dominates the approximation error and highlights the inadequacy of uniform time allocation. Experimental results support this analysis: as shown in Fig. 2, although the overall approximation error is small, the fitting degrades significantly in the near-zero region where outputs vary most rapidly. This region is crucial for Transformer performance, and the poor local approximation dominates overall behavior (FS-based Transformer conversion results in Appendix D).
IV Method
In this section, we first propose the MBE neuron, which employs exponential decay parameter update strategy and multi-basis encoding method to map activation values at multiple resolutions. Based on the MBE neuron, we further design an A2S conversion framework to overcome key challenges for Transformer architectures, achieving near-lossless conversion that is training-free and low-latency.
IV-A Multi-Basis Exponential Decay Neuron
To address the limited expressiveness of FS neurons, we propose a novel MBE neuron with an exponential decay parameter update strategy and multi-basis encoding scheme. Unlike FS neurons requiring learning multiple parameters $(d[t],r[t],V_{\mathrm{th}}[t])$ per timestep, MBE neurons only learn decay rates and discrete timesteps, reducing parameter overhead and gradient instability (Details in Appendix E).
Figure 3: FS neuron and MBE neuron encoding methods.
The exponential decay mechanism enables progressive refinement from coarse to fine resolution, allowing adaptive approximation of functions with varying granularity and rapid focus on limited value ranges over shorter timesteps, as illustrated in Fig. 3. The parameter update is defined as:
$Para(\tau_{n},t)=\alpha\cdot exp\left(-\frac{t\Delta t}{\tau_{n}}\right),$ (4)
where $Para(\tau_{n},t)$ represents the parameter value at time $t$, $\alpha$ is a hyperparameter typically set to the target function’s maximum value, $\Delta t$ is the discrete timestep, and $\tau_{n}\in\{\tau_{d_{n}},\tau_{r_{n}},\tau_{V_{{th}_{n}}}\}$ are the corresponding decay rates.
Figure 4: Multi-basis exponential decay neuron.
To further enhance the representational capacity for approximating diverse nonlinear functions, we introduce multi-basis encoding as the second core feature. As shown in Fig.4, each MBE neuron comprises $n$ basis components contributing distinct functional components to the overall response. The dynamics of the $n$-th basis are defined as:
$u_{n}[t+1]$ $=u_{n}[t]-s_{n}[t]\cdot r_{n}[t],$ (5)
$s_{n}[t]$ $=\mathcal{H}\big{(}u_{n}[t]-{V_{th}}_{n}[t]\big{)},$ (6)
$o_{n}[t+1]$ $=o_{n}[t]+s_{n}[t]\cdot d_{n}[t],$ (7)
where $u_{n}[t]$ is the membrane potential of basis $n$ at time $t$, $o_{n}[t]$ is the accumulated output, $d_{n}[t]$ denotes the spike intensity, $r_{n}[t]$ is the reset value, ${V_{th}}_{n}[t]$ is the firing threshold, and $\mathcal{H}(\cdot)$ is the Heaviside step function. When $u_{n}[t]\geq{V_{th}}_{n}[t]$, the neuron fires and emits a spike with intensity $d_{n}[t]$ while reducing the membrane potential by $r_{n}[t]$.
Finally, all basis outputs are weighted by a $w$ to form $\hat{f}(x)$, passed to the next layer to approximate $f(x)$:
$\hat{f}(x)=\sum\nolimits_{n=1}^{N}w^{(n)}\cdot\text{o}^{(n)}(T)\approx f(x).$ (8)
Approximation Error Analysis Following the same analytical framework in Theorem 1, we derive error bounds for MBE neurons, accounting for the representational capacity:
Theorem 2 (MBE Neuron Error Bounds).
For a target activation function $f:[a,b]\rightarrow\mathbb{R}$ and MBE neuron output $\hat{f}_{N,T}^{(M)}(x)$ with $N$ basis components, $T$ timesteps trained on $M$ samples, the total approximation error satisfies:
$\varepsilon^{*}\!\leq\!O\Biggl{(}\underbrace{\sqrt{\frac{N\log(N)\log M}{M}}}_{\text{Empirical Gap}}\!+\!\!\underbrace{\frac{\mathcal{L}_{f}|y|_{max}}{NT}}_{\text{Parametric Gap}}\!\!+\!\underbrace{\frac{\|w\|_{1}|\alpha|\tau_{max}}{T\Delta t}}_{\text{Quantization Gap}}\Biggr{)},$
where $\varepsilon^{*}=\mathbb{E}[|f(x)-\hat{f}_{N,T}^{(M)}(x)|]$, $\mathcal{L}_{f}$ is the Lipschitz constant of $f$, $|y|_{max}$ is the maximum absolute value of the target function, $\|w\|_{1}$ is the $L_{1}$ norm of basis weights, $|\alpha|$ is the positive scaling parameter, $\tau_{max}=\max_{n}\{\tau_{d_{n}},\tau_{r_{n}},\tau_{V_{th_{n}}}\}$ is the maximum time constant, and $\Delta t$ is the discrete timestep. The proof is provided in Appendix B.
The theoretical analysis shows MBE advantages and guides our conversion framework through three key insights:
Basis component design: The Parametric Gap $\mathcal{O}(1/NT)$ outperforms FS neurons’ $\mathcal{O}(\mathcal{L}_{f}|y|_{\max}/T)$ by enabling multi-basis encoding to adaptively focus on high-curvature regions, solving the GSO problem without extra timesteps.
Initialization stability: The Quantization Gap $\frac{\|w\|_{1}|\alpha|\tau_{max}}{T\Delta t}$ mitigates the EDI problem through controlled scaling via $|\alpha|$ and time constant effects via $\tau_{max}/\Delta t$, enabling initialization-stability parameter updates.
Hyperparameter determination: The $1/NT$ scaling reveals a principled trade-off: increasing basis components $N$ can effectively compensate for limited timesteps $T$, thereby achieving a balance between accuracy and low latency.
Figure 5: Overview of our framework, which depicts the approximations of FP multiplication, Softmax, LayerNorm, and GELU.
IV-B Conversion Framework
Based on MBE neurons, we propose a framework for Transformer-to-SNN conversion. As shown in Fig. 5, by decomposing non-spiking components (nonlinear activations, FP multiplication, Softmax, LayerNorm) into basic functions and designing corresponding MBE neurons for equivalent approximation, we enable spiking-based representation.
Approximation of Nonlinear Activation Functions
To enable accurate and efficient spike-based approximation of nonlinear activation functions such as GELU and Tanh in Transformers, we employ MBE neurons with $N=4$ basis components (in Eq.(8)). Since LayerNorm compresses activations into a narrow range, we constrain the GELU input domain to (120,10)(-120,10) to enhance both approximation accuracy and training stability. Within this interval, we uniformly sample M=10, [^0] M=10{,}000 points from the target function $f(x)$ and compute the spike-based output $\hat{f}(x)$ over $T$ timesteps.
Approximation of Floating-Point Multiplication
The FP multiplication operands $x_{1}$ and $x_{2}$ are transformed through approximate identity mapping by the MBE neuron $MBE_{Id}$, yielding the spike-train representations:
$x\approx MBE_{Id}(x)=\sum_{t=0}^{T-1}d[t]s[t],\quad x\in\{x_{1},x_{2}\}.$ (9)
Then, the multiplication of $x_{1}$ and $x_{2}$ is expressed as:
$x_{1}\cdot x_{2}$ $\approx\sum_{i=0}^{T-1}\sum_{j=0}^{T-1}d[i]d^{\prime}[j]\cdot s[i]s^{\prime}[j].$ (10)
We represent the spike sequences of two MBE neurons as vectors $\mathbf{s}=\{s[t]\}_{t=0}^{T-1}$ and $\mathbf{s^{\prime}}=\{s^{\prime}[t]\}_{t=0}^{T-1}$ respectively, and denote their corresponding intensity sequences as vectors $\mathbf{d}=\{d[t]\}_{t=0}^{T-1}$ and $\mathbf{d^{\prime}}=\{d^{\prime}[t]\}_{t=0}^{T-1}$. Based on these temporal vectors, we construct the intensity matrix $\mathbf{D}\in\mathbb{R}^{T\times T}$ and the spike matrix $\mathbf{S}\in\{0,1\}^{T\times T}$ as follows:
$\mathbf{D}$ $=\mathbf{d}^{\top}\mathbf{d^{\prime}},\qquad\mathbf{S}=\mathbf{s}^{\top}\mathbf{s^{\prime}}.$ (11)
To achieve multiplication in the form of spikes, we employ the Hadamard product followed by summation. Consequently, Eq.(10) can be expressed as:
$x_{1}\cdot x_{2}\approx\sum_{i=0}^{T-1}\sum_{j=0}^{T-1}\mathbf{D}_{ij}\cdot\mathbf{S}_{ij}=\sum_{i=0}^{T-1}\sum_{j=0}^{T-1}(\mathbf{D}\odot\mathbf{S})_{ij},$ (12)
where $\mathbf{D}_{ij}$ = $d[i]d^{\prime}[j]$ and $\mathbf{S}_{ij}$ = $s[i]s^{\prime}[j]$. The intensity matrix $\mathbf{D}$ can be precomputed and shared across the entire network. With the binary matrix $\mathbf{S}$, the entire FP multiplication approximation process maintains its spike-driven characteristics (Details in Appendix. F.1). By replacing FP multiplications in self-attention with our spike-driven operations, we enable fully spike-based attention matrix computation.
Approximation of Softmax
The $\text{Softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$, where $i$ ,$j$ denote the indices of elements in the sequence. Since softmax depends on all inputs, it cannot be directly implemented by a single MBE neuron. As shown in Fig. 5(b), we decompose the softmax computation into three components: exponential ($e^{x}$), reciprocal ($1/x$), and FP multiplication, each approximated by MBE neurons.
For the term $e^{x_{i}}$, we apply the change-of-base formula to decompose it into integer and decimal components as:
$e^{x_{i}}=2^{x_{i}\cdot\log_{2}e}=2^{\lfloor{x_{i}}\cdot\log_{2}e\rfloor}\cdot 2^{x_{i}-\lfloor{x_{i}}\cdot\log_{2}e\rfloor},$ (13)
where $\lfloor{\cdot}\rfloor$ denotes the floor operation. The decimal component $2^{x_{i}-\lfloor x_{i}\cdot\log_{2}e\rfloor}$ is approximated by MBE neurons, and the integer exponent $2^{\lfloor x_{i}\cdot\log_{2}e\rfloor}$ is achieved by hardware-efficient addition [28]. For the reciprocal term $1/\sum_{j}e^{x_{j}}$, we follow the IEEE 754 standard [29] to extract the exponent $E$ and mantissa $M$ through the corresponding bits of $\sum_{j}e^{x_{j}}$ (i.e., $\sum_{j}e^{x_{j}}=M\cdot 2^{E}$). The reciprocal $1/M$ is directly approximated using an MBE neuron, and the final result is obtained as $1/\sum_{j}e^{x_{j}}=2^{-E}/M$.
Combining the approximations of $e^{x_{i}}$ and $1/\sum_{j}e^{x_{j}}$, together with our FP multiplication approximation scheme, we compute the final softmax output in a spike-based manner (Detailed Softmax algorithm table in Appendix F.2).
Approximation of LayerNorm
The $\text{LN}(x_{i})=\gamma\cdot\frac{x_{i}-\mu}{\sqrt{\sum(x_{i}-\mu)^{2}/n+\epsilon}}+\beta$, where $\mu$ is the input mean, $n$ is the number of elements, $\epsilon$ is a stability constant, and $\gamma,\beta$ are learnable parameters. As shown in Fig. 5(c), we also decompose LayerNorm into three spike-unfriendly operations: squaring ($x^{2}$), reciprocal square root ($1/\sqrt{x}$), and FP multiplication, each approximated by MBE neurons.
For the squared term $\sum(x_{i}-\mu)^{2}$, we adopt the proposed FP multiplication approximation with the intensity matrix $\mathbf{D}$ pre-scaled by $1/n$ to directly yield $(x_{i}-\mu)^{2}/n$.
For the inverse square root $1/\sqrt{\sum(x_{i}-\mu)^{2}/n+\epsilon}$, we decompose the variance term into exponent $E$ and mantissa $M$. We adjust $E$ and $M$ based on $E$’s parity to ensure E/2-E/2 is an integer. The term $1/\sqrt{M}$ is approximated directly using MBE neuron, and the inverse square root is obtained by multiplying with $2^{-E/2}$, similar to the Softmax approximation. Finally, we multiply $(x_{i}-\mu)$ with the approximated $1/\sqrt{\sum(x_{i}-\mu)^{2}/n+\epsilon}$ using our FP multiplication approximation to obtain the LayerNorm output.
Based on the aforementioned conversion of all non-spike-friendly components into equivalent spiking forms, we replace the corresponding modules to construct the Transformer-based SNN without modifying the original network parameters. Detailed pseudocode is in Appendix F.3.
TABLE I: ImageNet Performance Comparison.
Category Method Arch. Spike-Driven Training-Free Param[M] Timesteps ANN Acc.[%]($\Delta$)
DT DSR[30] ResNet-18 - 12 50 - 67.74
TET[31] SEW-ResNet-34 - 22 4 - 68.00
AT-SNN[32] ResNet-104 - 45 4 - 77.08
Spikingformer[33] -4-384-400E - 66 4 - 75.85
SDTv1[34] -8-768 - 66 4 - 77.07
Spikeformer[35] -7L/3×2×4 - 38 4 - 78.31
QKFormer[36] HST-10-768 - 65 4 - 84.22
SDTv3[8] E-SpikeFormer - 173 8 - 85.10
A2S QCFS [16] VGG-16 138 64 74.92 72.85(-2.07)
QFFS[37] VGG-16 138 8 73.08 73.10(+0.02)
SNM[25] ResNet-18 12 64 73.18 71.50(-1.68)
QCFS[16] ResNet-34 22 64 74.23 72.35(-1.88)
AdaFire[38] ResNet-34 22 8 75.66 72.96(-2.70)
ECL[39] ResNet-34 22 16 74.36 72.37(-1.99)
Ours VGG16 138 10 73.37 72.61 (-0.76)
ResNet-34 22 10 76.31 75.57 (-0.74)
ECMT[27] ViT-L/16 307 12 84.88 84.71(-0.17)
STA[19] ViT-B/32 86 256 83.60 82.79(-0.81)
SpikeZIP-TF[18] SViT-L-32Level 304 64 85.41 83.82(-1.59)
SpikedAttention[26] Swin-T(BN) 28 48 79.30 77.20(-2.10)
Ours ViT-B/16 86 16 83.44 83.00(-0.44)
ViT-M/16 64 16 85.95 85.31(-0.64)
V Experiments
V-A Experimental Setup
We conduct experiments on a GPU (RTX 4090) environment using the PyTorch framework, evaluating both pre-trained Vision Transformer (ViT) models, including ViT-Base-Patch16 (ViT-B/16) [40] and ViT-Medium-Patch16-Reg4-Gap-256 (ViT-M/16) [41], as well as CNN models, including VGG16 [42] and ResNet34 [43] on the ImageNet dataset [44]. To validate the generalizability of our method across different language domains and tasks, we adapt RoBERTa [45] for natural language understanding (NLU) tasks and employ GPT-2 [46] for natural language generation (NLG) task evaluation (Experimental and implementation details in Appendix G.1).
V-B Comparative Study
Comparison on CV
We evaluate the performance of our conversion framework by applying it to convert both ViT and CNN architectures. As shown in Tab. I, ViT-B/16 and ViT-M/16 achieve conversion losses of 0.44% and 0.64% respectively—significantly lower than most existing methods. The framework attains 85.31% accuracy for ViT and 75.57% for CNN, outperforming other SNNs. While ECMT achieves competitive accuracy at shorter timesteps, it retains FP multiplication in its expectation compensation module. In contrast, our A2S framework maintains crucial spike-based property, thereby avoiding the high energy consumption typically associated FP multiplication during inference. Furthermore, the training-free nature of our approach enables conversion with minimal computational overhead, eliminating the need for extensive retraining for source ANNs. Our method successfully accomplishes A2S conversion without requiring additional training of the original network, preserves the essential spiking properties, delivers superior performance at short timesteps, and achieves optimal results across all evaluation dimensions.
Comparison on NLU
To evaluate our method on NLU tasks, we conduct experiments on four benchmark datasets including SST-2, SST-5, MR, and Subj, following standard experimental settings from studies [18, 47]. As shown in Table II, our method achieves competitive performance across all datasets, outperforming existing methods on both RoBERTa-Base (125M) and RoBERTa-Large (355M). On SST-2, it attains 95.98% accuracy with $T=16$, representing only 0.24% degradation from the original ANN (96.22%), while outperforming SpikeZIP-TF [18] by 2.19% and reducing timesteps by 87.5%. Similarly, on MR with RoBERTa-Base, our method achieves 89.00% accuracy at $T=16$, exceeding SpikeZIP-TF by 2.87% with 75% fewer timesteps.
Comparison on CV
We evaluate the performance of our Spike-F3 framework by applying it to convert both ViT-B/16 and ViT-M/16 architectures. The experiment results demonstrate that our Spike-F3 method achieves SOTA performance across multiple metrics. As illustrated in Tab. I, the conversion loss for the ViT-B/16 and ViT-M/16 models is measured at 0. 44% and 0. 64%, respectively, significantly lower than most existing methods, while achieving high accuracy 85.31% at T=16. Although ECMT[27] achieves good accuracy at shorter timesteps, it keeps FP multiplication in the expectation compensation module for matrix product. In contrast, converted SNNs based on Spike-$\text{F}^{3}$ maintain crucial spike-based property, thereby avoiding the high energy consumption typically associated FP multiplication during inference. Furthermore, the training-free nature of our approach enables conversion with minimal computational overhead, eliminating the need extensive retraining for source ANNs. Our method successfully accomplishes nearly lossless Transformer A2S conversion while avoiding retraining of the original ANN. It effectively preserves essential spiking neural dynamics, demonstrates superior inference accuracy under low timesteps conditions, and ultimately achieves SOTA performance across all comprehensive evaluation metrics.
TABLE II: NLU Performance Comparison
Category Model Param SST-2 SST-5 MR Subj T
ANN Roberta[45] 125 94.49 55.46 89.39 96.45 1
355 96.22 59.37 91.36 97.50 1
DT Spikeformer[35] 110 81.55 42.02 79.38 91.80 4
SpikeBERT[48] 109 85.39 46.11 80.69 93.00 4
SpikeGPT[47] 45 80.39 37.69 69.23 88.45 50
A2S SpikeZIP-TF[18] 125 92.81 52.71 86.13 95.55 64
355 93.79 56.51 89.28 96.70 128
Ours 125 93.46 55.11 89.00 96.30 16
355 95.98 58.31 90.96 97.45 16
Comparison on NLU
To assess the effectiveness of Spike-F3 in NLU tasks, we conduct experiments on four benchmark datasets including SST-2, SST-5, MR, and Subj, following standard experimental settings from previous studies [18, 47]. As shown in Tab. II, Spike-F3 achieves SOTA performance across all evaluated datasets, consistently outperforming existing methods on both RoBERTa-Base (125M) and RoBERTa-Large (355M). On SST-2, it achieves 95.98% accuracy with T=16, resulting in only 0.24% conversion loss compared to the original ANN (96.22%). Compared to SpikeZIP-TF [18], Spike-F3 improves accuracy by 2.19% and reduces timesteps by 87.5%. In the lightweight setting, Spike-F3 achieves 89.00% accuracy on MR using RoBERTa-Base with T=16, surpassing SpikeZIP-TF by 2.87% while reducing timesteps by 75%. These results confirm the effectiveness of Spike-F3 in training-free A2S for NLU tasks.
TABLE III: NLG Performance Comparison. Both Wiki-2 and Wiki-103 use perplexity. indicates lower is better.
Category Model Param [M] Wiki-2 $\downarrow$ Wiki-103 $\downarrow$ T
ANN GPT-2[46] 346 22.34 22.65 1
DT SpikeGPT[47] 216 18.01 39.75 1024
A2S Spike-F3 (Ours) 346 22.69 (+0.35) 23.41 (+0.76) 16
Figure 6: (a) Effect of MBE and FS neuron on components in Transformer, FP in (a) means floating multiplication. (b) Impact of MBE neurons with varying basis counts on energy consumption and accuracy when replacing GELU. MR represents marginal rate of change between accuracy and energy consumption. (c) Effect of the basis and decay. Loss comparison between models with (blue) and without (red) decay mechanism across basis numbers N=1,2,4,6,8. Y-axis shows log MSE loss (lower is better).
Comparison on NLG
To demonstrate the applicability of our method to NLG tasks, we evaluate its performance on WikiText-2 and WikiText-103. As shown in Tab. III, our method achieves 22.69 perplexity on WikiText-2 with only T=16 timesteps, incurring merely 0.35% conversion loss compared to GPT-2. More significantly, on WikiText-103, it attains 23.41 perplexity—a substantial 41.1% improvement over directly-trained SpikeGPT (39.75) while using dramatically fewer timesteps (T=16 vs. T=1024), closely approaching the original GPT-2 performance and demonstrating superior efficiency for long-context generation.
V-C Ablation Study
To compare MBE neurons and FS neurons, we evaluate their performance on key Transformer components. For GELU and FP multiplication, direct replacement is employed. As shown in Fig. 6(a), MBE neurons surpass FS neurons by $13.20\%$ in GELU conversion and achieve improvements in Softmax and LayerNorm, demonstrating superior approximation ability. In Fig. 6(b), increasing $N$ leads to more accurate function approximations by MBE neurons. MBE neurons achieve high performance (e.g., less than 1% conversion loss when N $\geq 4$) while offering significantly higher energy efficiency than ANNs. N=4 achieves the optimal accuracy-MR trade-off, representing the knee point where marginal gains plateau. The synergistic effect between exponential decay and multi-basis encoding is demonstrated in Fig. 6(c). Incorporating decay consistently reduces MSE by 1122 orders of magnitude across all basis numbers, with significant performance improvements. This reveals that exponential decay creates a multiplicative synergy with multi-basis encoding, enabling exceptional approximation precision unattainable through basis expansion alone.
TABLE IV: Performance across various timesteps and models. For Wiki-103, perplexity () is used; lower is better.
Dataset Model ANN Timestep (T)
8 10 12 16
ImageNet ViT-B/16 83.44 0.12 79.96 82.79 83.00
ViT-M/16 85.95 1.17 21.99 84.43 85.31
MR Roberta-B 89.39 50.17 71.30 88.55 89.00
Roberta-L 91.36 49.66 68.62 90.52 90.96
Wiki-103 GPT-2 22.65 41072 11992 33.56 23.41
We evaluate ImageNet, MR, and Wiki-103 to analyze timestep requirements for A2S conversion across vision and language tasks. As shown in Tab. IV, our method achieves about 1% conversion loss for ViT, RoBERTa, and GPT-2 at T=16, with near-optimal performance at T=12, demonstrating efficient progressive encoding. Notably, ViT-M/16 shows a 21.99% accuracy drop at T=10 vs. ViT-B/16, primarily due to its wider identity mapping range [0,62] causing amplified approximation errors under limited timesteps.
VI Energy Estimation
Unlike ANNs where energy depends on floating-point operations (FLOPs), the energy cost of SNNs is dominated by synaptic operations (SOPs). Following [49], the energy ratio between SNNs and ANNs is:
$\frac{E_{\text{SNN}}}{E_{\text{ANN}}}=\frac{\text{SOPs}\cdot E_{\text{AC}}}{\text{FLOPs}\cdot E_{\text{MAC}}},$ (14)
where $E_{\text{MAC}}$ = 4.6pJ, $E_{\text{AC}}$ = 0.9pJ. For ViT-M/16, we measure the firing rates $\eta$ of nonlinear operations across all layers. Taking GELU as an example, the ANN implementation requires 70 FLOPs, while the MBE neuron achieves $\eta$=38.22% under N=4 and T=16. The synaptic operations achieve energy consumption of 13.7% according to $T*N*\eta*E_{AC}$. Other operations yield greater energy savings owing to lower MBE firing rates. Complete results and detailed firing rates are provided in Appendix G.4.
VII Conclusion
This paper proposes an efficient Spiking Transformers conversion method that requires no additional training on the source ANNs. We first theoretically and experimentally identify two key challenges in using FS neurons to approximate nonlinear operations: excessive dependence on initialization (EDI) and global suboptimality (GSO) problems. To address these issues, the MBE neuron employs an exponential decay strategy and a multi-basis encoding method to more accurately approximate the nonlinear operations in Transformer architecture. Based on the MBE neuron, a general ANN-to-SNN conversion framework is developed, which supports spiking nonlinear activation functions, spiking FP multiplications, spiking Softmax, and spiking LayerNorm. Extensive experiments on various models (CNN, ViT, RoBERTa, GPT-2) across tasks (CV, NLU, NLG) confirm that the proposed method achieves near-lossless conversion. Therefore, our method provides a promising approach for the efficient and scalable deployment of Transformer-based SNNs in real-world applications.
E.2 Approximation of Nonlinear Operations
Figure 8: Comparison of FS and MBE neurons in approximating various nonlinear functions.
To more clearly compare our method with FS neurons in approximating nonlinear operations, we evaluate a series of representative functions, including SiLU, $1/x$, $e^{x}$, $\sqrt{1/x}$ and GELU. For the SiLU function, we follow the same evaluation protocol as used in the FS paper, measuring approximation errors separately in key and non-key regions. The comparison results are summarized in Tab. VII, while the results for the other functions are shown in Fig. 8.
TABLE VII: Accuracy comparison between FS and MBE neuron in approximating the SiLU function across different intervals. Lower values indicate better approximation.
Interval [8,2][-8,\,-2] [2, 2][-2,\,2] [2, 12][2,\,12]
FS 0.0064 0.0023 0.0064
MBE (Ours) $6.97\times 10^{-5}$ 0.0006 0.0048
Improvement (-98.91%) (-73.91%) (-25.00%)
-F Conversion Implementation via MBE Neurons
TABLE VIII: Steps of Spike-based Implementation for Softmax
Step Operation Formulation Spike Implementation Description
0 Input $x_{i}\ (i\in[1,K])$ Raw input tokens Input sequence to be normalized
1 Exponential $e^{x_{i}}=2^{x_{i}\log_{2}e}$ $e^{x_{i}}\approx MBE_{\exp}(x_{i},T,B)$ Change-of-base decomposition
2 Summation $S=\sum_{j=1}^{N}e^{x_{j}}$ $S=\sum_{j=1}^{N}MBE_{\exp}(x_{j},T,B)$ Accumulate all exponential outputs
3 Reciprocal $\frac{1}{S}=\frac{1}{M}\cdot 2^{-E}$ $\frac{1}{S}\approx MBE_{\text{inv}}(M,T,B)\cdot 2^{-E}$ Extract IEEE 754 format
4 Multiplication $y_{i}=e^{x_{i}}\cdot\frac{1}{S}$ $y_{i}=\text{SpikeMul}(MBE_{\exp}(x_{i}),MBE_{\text{inv}}(S))$ Spike-based matrix multiplication
5 Output $\hat{y}_{i}=\text{Softmax}(x_{i})$ Spike-driven Softmax output Final normalized distribution
F.1 MBE Neuron for FP Multiplication
Initially, we perform random sampling on the dataset to statistically determine the value ranges of the two activation values involved in floating-point multiplication. As illustrated in the upper portion of Fig. 9, we use MBE neurons to approximate an identity mapping over this range and convert the involved floating-point values into spike-weighted summation forms. The approximate product of the weighted sums is given by:
$x_{1}\cdot x_{2}$ $\approx\sum_{i=0}^{T-1}d[i]s[i]\cdot\sum_{j=0}^{T-1}d^{\prime}[j]s^{\prime}[j],$ (22)
where $d[t]$ and $d^{\prime}[t]$ denote the spike intensities at time step $t$, and $s[t]$ and $s^{\prime}[t]$ represent the spike events at time step $t$. Since the intensity values are in floating-point format, directly multiplying the spike-weighted sums still entails floating-point arithmetic, which is not spike-friendly. To address this, we equivalently transform the multiplication expression to decouple the spike terms from the floating-point components. Specifically, we reformulate Eq.(22) as:
$x_{1}\cdot x_{2}$ $\approx\sum_{i=0}^{T-1}d[i]d^{\prime}[i]\cdot\sum_{j=0}^{T-1}s[j]s^{\prime}[j]$ (23)
$=\sum_{i=0}^{T-1}\sum_{j=0}^{T-1}d[i]d^{\prime}[j]\cdot s[i]s^{\prime}[j].$ (24)
Based on Eq. 24, the factors of the accumulation terms can be represented as distinct vectors, which are formally defined as follows:
$\mathbf{i}=[i[0],i[1],\dots,i[T-1]],\quad\text{where}\quad\mathbf{i}\in\{\mathbf{d},\mathbf{d^{\prime}},\mathbf{S},\mathbf{S^{\prime}}\},$ (25)
where $\mathbf{S}$ and $\mathbf{S^{\prime}}$ represent the temporal spike train of MBE neuron while $\mathbf{d}$ and $\mathbf{d^{\prime}}$ denote the corresponding spike intensity. As illustrated at the bottom of Fig. 9, the intensity matrix $\mathbf{D}$ is obtained through the outer product of two neuron-specific constant intensity vectors. Similarly, the binary spike matrix $\mathbf{S}$ is formed in an analogous manner and can be expressed as:
$\mathbf{D}$ $=\mathbf{d}^{\top}\mathbf{d^{\prime}},\qquad\mathbf{S}=\mathbf{s}^{\top}\mathbf{s^{\prime}}.$ (26)
Subsequently, we compute the element-wise Hadamard product of $\mathbf{D}$ and $\mathbf{S}$, followed by summation to yield the approximated floating-point multiplication result:
$x_{1}\cdot x_{2}\approx\sum_{i=0}^{T-1}\sum_{j=0}^{T-1}\mathbf{D}_{\text{ij}}\cdot\mathbf{S}_{\text{ij}}=\sum_{i=0}^{T-1}\sum_{j=0}^{T-1}(\mathbf{D}\odot\mathbf{S})_{ij}.$ (27)
Figure 9: Approximation of Floating-Point Multiplication Using MBE Neurons. Top: Direct approximation of floating-point multiplication. Bottom: Spike-based approximation reformulated as a Hadamard product.
The spike-driven nature of this process is ensured through two key mechanisms: intensity matrix construction and spike-based operation.
Intensity matrix construction
The weight matrix depends solely on the neuronal spike intensity $d[t]$, which is determined during training. Consequently, $\mathbf{D}$ becomes an input-independent quantity that can be precomputed and stored prior to inference.
Spike-based operation
The Hadamard product involves only floating-point numbers and spikes, eliminating any floating-point-to-floating-point operations.
By acquiring MBE neuron spike intensities, we construct intensity matrix $\mathbf{D}$, while spike matrix $\mathbf{S}$ is obtained through T-step AND operations in inference. Floating-point multiplication is approximated via efficient Hadamard product summation. When activation values originate directly from MBE neurons, Eq. 26 enables direct matrix construction without additional spike-generating neurons.
F.2 MBE Neuron for Softmax As depicted in Tab  VIII, the process begins with raw input tokens $x_{i}$ where $i\in[1,K]$ that undergo five key computational steps. First, the exponential operation $e^{x_{i}}$ is reformulated using the change-of-base formula $e^{x_{i}}=2^{x_{i}\log_{2}e}$ and approximated by MBE neurons specialized for exponential computation. Next, the summed exponential terms $S=\sum_{j=1}^{K}e^{x_{j}}$ are computed through iterative spike-based summation of individual outputs. The reciprocal operation $\frac{1}{S}$ is then implemented by decomposing it into IEEE 754 floating-point format components, leveraging MBE neurons for both mantissa and exponent $2^{-E}$ calculations. The final multiplication $y_{i}=e^{x_{i}}\cdot\frac{1}{S}$ is performed via spike-based matrix multiplication between the exponential and reciprocal MBE outputs. This complete pipeline culminates in spike-driven Softmax output $\hat{y}_{i}=\text{Softmax}(x_{i})$, achieving analog computation of the probability normalization function through discrete spike events while maintaining mathematical equivalence to the continuous Softmax formulation.
F.3 Overall Process Pseudocode See Algorithm 1.
Algorithm 1 Conversion Procedure of Proposed Framework
1:Input: Pre-trained Transformer ANN model $f_{\text{ANN}}(W)$; Dataset $D$; Timesteps $T$ and basis $B$ of MBE neurons.
2:Output: SNN model $f_{\text{SNN}}(W,T,B)$
3:Step 1: Collect intervals $R$ of MBE neurons
4: Sample minibatch data from $D$.
5: Run the data on $f_{\text{ANN}}$ and record input ranges of nonlinear operations in: FP multiplication, nonlinear activation functions, Softmax, LayerNorm.
6:Step 2: Construct Spiking Modules
7: Train MBE neurons with timestep $T$ and basis $B$ using collected intervals $R$.
8: Build spiking implementations of:
9:  FP multiplication
10:  nonlinear activation functions
11:  Softmax & LayerNorm
12:Step 3: Model Conversion
13:for each module $m$ in $f_{\text{ANN}}$ modules do
14:  if $m$ is Matrix Multiplication then
15:   Replace $m$ with spiking FP multiplication
16:  else if $m$ is Activation Function then
17:   Replace $m$ with spiking nonlinear activation
18:  else if $m$ is Softmax then
19:   Replace $m$ with spiking Softmax
20:  else if $m$ is LayerNorm then
21:   Replace $m$ with spiking LayerNorm
22:  end if
23:end for
24:return $f_{\text{SNN}}(W,T,B)$
-G Supplementary Experimental Materials
G.1 Implementation Details The number of basis $N$ and timesteps $T$ are determined according to the complexity of the target functions being approximated. For GELU and Tanh approximation, we employ MBE neurons configured with $N=4$ and $T=16$. To mitigate approximation errors in identity mapping, all learnable parameters in the MBE neuron are fixed as [20]. Using this configuration with $N=8$, $T=16$, we implement function approximations for three distinct mathematical operations: the exponential function $2^{x}$ over the interval [0, 1], the reciprocal function $1/x$ within [0.5, 1], and the inverse square root function $1/\sqrt{x}$ across [0.5, 2]. A unified training protocol is adopted, with a fixed learning rate of 0.01 for 200 epochs. To further enhance training stability, an exponential optimizer with a decay rate of 0.99 is used to ensure stable convergence. G.2 Evaluation of CNN architectures To further validate the broad applicability of the proposed framework, we evaluate it on two representative CNN architectures: VGG16 and ResNet34. As shown in Tab. IX, the framework incurs only a 0.75% accuracy drop at T=10, demonstrating its effectiveness and competitiveness compared to state-of-the-art (SOTA) methods.
TABLE IX: Performance comparison of different methods on VGG16 and Resnet34 architectures
Method Arch. ANN T Acc.[%]($\Delta$)
QCFS [16]
VGG16 74.29 32 68.47 (-5.82)
SRP [14]
VGG16 74.29 16 69.13 (-5.16)
SNM [25]
VGG16 73.18 64 71.50 (-1.68)
Ours VGG16 73.37 10 72.61 (-0.76)
QCFS [16]
Resnet34 74.23 64 72.35 (-1.88)
ECL [39]
Resnet34 74.36 16 72.37 (-1.99)
AdaFire [38]
Resnet34 75.66 8 72.96 (-2.70)
Ours Resnet34 76.31 10 75.57 (-0.74)
G.3 Ablation Study
TABLE X: Effect of N and decay in MBE neuron. Func represents ”Function” and * means MBE neurons without decay.
Func. N=1 N=2 N=4 N=6 N=8 N=8*
GELU 7.1e-03 4.1e-03 2.3e-04 1.7e-04 1.0e-04 2.8e-01
invsqrt 1.4e-03 1.2e-03 3.1e-04 1.0e-04 4.9e-05 2.8e-03
inv 8.6e-03 2.7e-03 1.1e-03 2.2e-03 4.4e-04 9.8e-03
$2^{x}$ 8.9e-04 4.5e-04 4.0e-04 2.4e-04 5.3e-05 4.5e-03
The proposed multi-basis encoding architecture employs a hierarchical set of specialized basis kernels, each designed to capture distinct value ranges, thereby enabling high-fidelity approximation through adaptive multi-scale representation. Concurrently, the exponential decay mechanism facilitates progressive refinement via iterative parameter attenuation, effectively mitigating premature saturation while preserving dynamic representation capacity.
To quantitatively evaluate the individual contributions of these components to approximation accuracy enhancement, we conducted comprehensive ablation studies. As demonstrated in Tab.  X, systematic expansion of the basis count (N=1→8) yields significant MSE reduction across all evaluated nonlinear operators (GELU, invsqrt, inv, and $2^{x}$), with invsqrt achieving a remarkable 96.46% improvement. Crucially, ablation of the decay mechanism (N=8 (w/o decay)) results in substantial performance deterioration, unequivocally establishing its fundamental role in ensuring approximation stability and precision.
G.4 Overall Energy Estimation
TABLE XI: Firing rates of MBE neuron when fitting various functions at T=16.
Operations FiringRate(%)
Attention_score 8.31
$2^{x}$ 46.94
$1/x$ 3.74
$1/\sqrt{x}$ 25.27
LayerNorm_input_identity 27.61
LayerNorm_1/x_identity 38.46
GELU 38.22
The energy consumption of a single MBE in $T$ time steps is calculated as $E_{MBE}=T\times\eta\times N\times C\times N_{h}\times EAC$, where $T$ is the timestep, $\eta$ is the average firing rate, $N$ is the number of tokens, $C$ is the number of token channels, and $N_{h}$ is the number of heads (1 if it is a non-Attention operation). The energy consumption calculation formula for floating-point multiplication is $E_{FP}=T^{2}\times\eta_{1}\times\eta_{2}\times MO\times EAC$, where $\eta_{i}$ is the average firing rate of the MBE neuron $i$ of the spike source, and $MO$ is the number of matrix multiplication operations. Following the energy estimation model, the normalized energy ratio between an SNN module and its ANN equivalent is defined as:
$\frac{E_{\text{SNN}}}{E_{\text{ANN}}}=\frac{\text{SOPs}\cdot E_{\text{AC}}}{\text{FLOPs}\cdot E_{\text{MAC}}},\quad E_{\text{MAC}}=4.6\,\text{pJ},E_{\text{AC}}=0.9\,\text{pJ}.$ (28)
For the ViT-M/16 model, we analyze the spike firing rates of MBE neurons across various components during inference at $T=16$ via random sampling. Subsequently, as summarized in Tab. XI, we calculate the operation counts for each component within a single block. Based on the average energy consumption per block, the energy ratio of the converted SNN compared to the original ANN is 17.29 % according to Eq.(28). The result illustrates that our framework encodes information efficiently with sparse spikes while retaining most of the original ANN accuracy.